# Contrastive Localized Language-Image Pre-Training

### v1

@@ 6 Conclusion
Please see Appendix C for more discussions where we comment on the limitations, future directions, computation cost, design rationales, etc.
We tackle a deficiency of CLIP, to make the semantics aligned in the vision space not only at the image level but also at the region level. We propose a new pre-training framework that includes innovations in a new learning formulation that a strong encoder should be easily transformed in the foresee of downstream use of MLLMs. Our encoder creates a new possibility for adapting the features with input prompts of interaction together with MLLMs. To resolve the need for large-scale region-text training data, we carefully design a pseudo-labeling pipeline for visually-enriched and spatially-localized captions. Our pre-trained encoder is essentially a drop-in replacement of CLIP, with competitive image-text performance, and extra capability demonstrated in region-text tasks and VQA tasks with MLLMs.
Acknowledgement
@@
Reproducibility Statement
We made our best efforts to exhaustively state the implementation details. Training hyper-parameters and model architectures are discussed in Section 3.23.3, and 5.1, with a summary in Appendix A and Table A. For evaluation, as mentioned in Section 5.1, we strictly follow the official setup with the codebase released by the original authors if applicable, with details provided in Section A.2. For our datasets, we provide data processing details in Section 4 and example codes in Appendix B. We are working hard on releasing the annotations with internal approvals.
@@
Appendix A Experiment Details
Below, we provide more details on pre-training hyper-parameters and downstream evaluation tasks.
A.1 Pre-training Hyper-parameters
For pre-training both the in-house CLIP baseline and CLOC, we mainly follow the hyper-parameters in Radford et al. (2021) with our in-house pre-training datasets. The images used for model training are identical for CLIP and CLOC, while CLOC is trained on the extra region-text annotations of the same images via the proposed VESL pipeline (details in Section B). Table A summarizes the training hyper-parameters used for all experiments and the setup for components specific to CLOC.
@@ A.1 Pre-training Hyper-parameters
A.2 Evaluation Tasks
We provide more details on the tasks constructed for evaluating the encoders in Section 5.
@@ A.2 Evaluation Tasks
Our CLOC training augments a new capability for CLIP to generate region-level embeddings. This enables us to perform zero-shot region-text tasks, in analogy to the image-text zero-shot tasks like ImageNet classification and COCO image-text retrieval that CLIP has been evaluated on.
@@
For region retrieval, similarly, the CLOC model encodes both the image regions and the region captions from the public region-text GRIT dataset that the regions are annotated by the Kosmos-2 pipeline (Peng et al., 2023). We randomly sampled a $2$K image validation set for fast evaluation. We have verified it is statistically stable compared to the whole set that contains about $20$M in total. Unlike image-text retrieval where the image captions are likely unique, the objects in regions of many images might be duplicated. Therefore, we opt to report recall@10 rather than recall@1 for GRIT region retrieval in Table 2.
@@
To demonstrate our CLOC can benefit MLLM downstream tasks as a better image backbone, we consider two sets of MLLM experiments.
First, we experiment with Ferret (You et al., 2023) that is capable of taking spatial referring inputs for grounding and referring VQA tasks. Ferret can consume a point, a bounding box, or a free-form referring. It designs a complicated visual sampler module that involves point sampling and $k$NN grouping. We suggest the readers refer to Figure 3 and Section 3.2 in You et al. (2023) for more details. Here, we consider two methods to use CLOC for Ferret model training: ($i$) we only take the ViT encoder in CLOC to replace the CLIP ViT and still use the original Ferret visual sampler; ($ii$) we further replace the visual sampler with our simple Prompter (essentially a lightweight transformer encoder with box positional encodings) in Section 3.3 as illustrated in Figure 1(3b). More specifically, we simply convert all types of spatial referring as boxes. As evidenced by Table 3 and Table 4, our Prompter can indeed be a much simpler alternative and may perform even better as it is more consistent with CLOC pre-training.
Second, we evaluate on general VQA tasks that do not consider extra spatial referring inputs. The pre-trained ViT of CLOC is a drop-in-replacement of CLIP ViT in two sets of experiments of LLaVA-1.5 (Liu et al., 2023) and LLaVA-NeXT (Liu et al., 2024). The main difference includes different supervised fine-tuning (SFT) datasets. Also, LLaVA-NeXT uses the AnyRes technique that decomposes an image into several subimages that are encoded independently with the ViT and concatenated together as the input for the decoder. LLaVA-1.5 by default freezes the ViT, while LLaVA-NeXT fine-tunes all parameters during SFT. Since the official LLaVA-NeXT is trained on some proprietary datasets that are not reproducible, we use the Open-LLaVA-NeXT repository. [^8] https://github.com/xiaoachen98/Open-LLaVA-NeXT Our experiments in Table 5 demonstrate CLOC not only improves general VQA tasks besides referring and grounding tasks, but also generalizes well for both LLaVA-1.5 and LLaVA-NeXT settings.
Appendix B VESL Data Engine
We provide more details about our pseudo-labeling data pipeline proposed in Section 4.
B.1 Implementation Details
@@
For re-captioning, the goal is to replace alt-text with long, diverse, and detailed captions that can be used to generate more visual concepts as the region candidate phrases for the OV detector. Technically, any strong image captioner can be an option. In our paper, we adopt the VeCap pipeline (Lai et al., 2024) and leverage their images with enriched captions.
To extract region phrase candidates from the long captions, we adopt name entity recognition (NER) to extract leaf entities from the captions, inspired by Zhang et al. (2022). The code listing below shows the Python example implementation, where stop words and common generic words are filtered, following Minderer et al. (2024).
Generating bounding boxes and assigning region captions can be done by querying an OV objection detector. We adopted the OWLv2 detector (Minderer et al., 2024) with their pre-trained L/14 checkpoint to annotate inputs with $448\times 448$ image resolutions.
@@
Listing 1: Python example code for Step 2 of VESL in Section 4 for extracting text candidate queries from a caption.
B.2 More Visualizations
As mentioned in the remarks of Section 4, we found the alt-text sourced from the original web-crawled images might not contain enough details describing the regional content, thus limiting the diversity and quality of the text candidate queries for the OV detector to detect more meaningful objects. In Figure A, we show some cherry-picked examples (since the web-crawled images are quite noisy) to demonstrate the reasons why high-quality captions can help our region-text annotation pipeline. In Minderer et al. (2024), the queries are generated by the $n$-grams of the alt-text, while ours are by NER as described in Section B.1 on top of the visually-enriched captions. Note that, in both methods we use the same OV detector but with different approaches to generate the queries.
As shown in Figure A, for simple images like in the first row, both methods are doing reasonably well to detect “message card”. However, when the scene becomes complicated (e.g., the second row), our methods can detect more objects since more visual concepts can be extracted from our rich caption as queries for the detector. Similarly, it can be seen that our method captures more items that the alt-text missed, e.g., “banana”, “eggs”, “butter”, etc. in the third row; “drawstring” in the fourth row; “apples” and “vases” in the last row. Also, it is more likely to extract a more detailed description of the region rather than a class name, such as “green-roofed cottage nestles” in Figure 3 and “decorative metal tree sculptures” for the image in the last row of Figure A. We believe such high-quality region labels essentially contribute to better supervision for CLOC pre-training.
Figure A: Examples comparing our VESL and the labeling approach in Minderer et al. (2024) that directly uses the $n$-grams of the crawled alt-text. For VESL, each image is annotated with the visual-enriched caption to replace the alt-text, which is used to generate region text candidates that capture the image content better.
Appendix C More Discussions
@@
One limitation for CLOC is the labeling efforts in preparing the training data. As we discussed in Section 1, there are no public large-scale region-text datasets since it is expensive to infer such labels up to the scales we consider here. Unlike previous work (Zhong et al., 2022) that cropping boxes from images for annotating, our VESL pipeline conducts inference in image-level, thus the cost does not scale with the number of detected regions. With that being said, such inference still requires hundreds of GPUs running in parallel for days to scale up to billions of images. We are working on releasing the annotations to accelerate future research for the community.
For CLOC, we focus on the training objective and framework formulation, while making minimal efforts on hyper-parameter tuning, architecture search, dataset cleaning, and etc., thus better performance could be achieved. Besides, although we have included extensive standard evaluation tasks, the fine-grained region knowledge could also be useful on more under-explored tasks.
@@
We suggest promising future directions. In Section 3.2, our Prompter formulation can take flexible prompts to guide the embeddings for specific tasks. In this work, we consider a prompt as a single bounding box or a text caption, but it has the potential to expand to various types such as points, a mask, users’ free-form referring, or multiple prompts in multiple types together. We believe a more versatile Prompter with co-designs for different objectives can have a big potential. Similarly, our VESL labeling pipeline limits to the detection box format. Annotators supported for more formats may further boost it. We believe our approach is promising, as more attention has been drawn recently for better image re-captioning (Li et al., 2024; Fan et al., 2024) that VESL relies on. In addition, CLOC provides a new capability to extract region features without further training, thus can be used as a foundation model for exploring new VL applications.
@@
We comment on the computation cost of our framework. Our large models (ViT L/14) were trained on $1024$ v5p TPUs for about $6$ days. To optimize Equation 2, CLOC needs extra computation. The main overheads come from the contrastive matrix but not the lightweight Prompter. Fortunately, we found it feasible since ($i$) only a few boxes in each image need to be sampled per update; ($ii$) the loss computation becomes a smaller proportion when the ViT scales up. Overall, we found the computation acceptable compared to CLIP. More memory-efficient optimization like SigLIP (Zhai et al., 2023) can be implemented with JAX shard_map [^9] https://jax.readthedocs.io/en/latest/jep/14273-shard-map.html ops.
Discussions on design rationales.
@@
(1) Why not use other locality-enhanced image encoders? We would like to note that many image encoders with localization capabilities like DINOv2 (Oquab et al., 2023), OWLv2 (Minderer et al., 2024), CLIP-Self (Wu et al., 2023) etc. are developed specifically for dense vision tasks that cannot perform zero-shot image-level tasks like CLIP and CLOC. Our goal is to develop a drop-in replacement for the CLIP encoder that improves localization while maintaining all of CLIP’s original strengths, such as zero-shot image-level tasks and its critical role as the image backbone for MLLMs. Furthermore, perhaps well-known within the MLLM community, these encoders have been shown in recent reports that they are not comparable enough to compete with CLIP as the vision backbone for MLLM tasks (Tong et al., 2024) due to CLIP’s superiority in vision-language alignment. We thus believe enhancing CLIP itself is more demanding as this paper focuses on.
@@
Although it sounds like a plausible approach, we would like to point out that contrastive pre-training and object detection are fundamentally different in their technical rationales. CLIP pre-training is often conducted on large batches of low-resolution and noisy images, while object detection is trained on small batches of high-resolution images. CLIP is by default trained from scratch and object detection is typically initialized from pre-trained encoders and focuses on the detection head. Furthermore, detection requires heavy computation on box proposals to detect all boxes appearing in an image, while our region-text contrastive design allows us to flexibly sample fewer regions per image as motivated in Equation 3. Overall, their data pipeline and distributed training setup are not on the same scale thus such joint training may not be very reasonable.
With that being said, some previous works do have attempts that are the exceptions but only for some but not all of the mentioned aspects, and mainly for the purpose of detection. For instance, DetCLIP-v2 (Yao et al., 2023) adds image-text contrastive loss into detection loss to improve open-vocabulary capability for detection. OWLv2 pre-trains the detector with rather small resolutions but still with a batch size of a maximum $256$ since each image will need to predict up to $100$ boxes during training. Both DetCLIP-v2 and OWLv2 are fine-tuned from a pre-trained encoder.
On the contrary, we study pre-training the image encoder from scratch, which may be complementary to the previous efforts. CLOC maximizes the similarity in co-design with CLIP, thus making it much easier to develop within the same codebase.
(3) Do we really need to train CLOC from scratch? What if we fine-tune from CLIP? As CLIP pre-training is expensive, one may wonder if it is necessary to train from scratch on the proposed region-text datasets, or if we can initialize from a standard CLIP trained on image-text pairs only and fine-tunes with CLOC for a shorter stage. Our early investigation, even with extensive hyper-parameter tuning, suggests it is likely to be suboptimal compared to training from scratch directly. For instance, we initialize from the CLIP model 2 in Table 2 and fine-tunes it for another extra $100$K steps with the CLOC training loss Equation 4. The model reaches $64.1\%/19.1\%$ mAcc on COCO/LVIS region recognition, which is much worse than $70.1\%/27.2\%$ of the model 8 trained from scratch, even with more overall training steps.

### v2

@@ 6 Conclusion
Please see Appendix D for more discussions where we comment on the limitations, future directions, computation cost, design rationales, etc.
We tackle a deficiency of CLIP, to make the semantics aligned in the vision space for both image and region level. We propose a new pre-training framework that includes innovations in a new learning formulation that a strong encoder should be easily transformed in the foresee of downstream use of MLLMs. Our encoder creates a new possibility for adapting the features with input prompts of interaction together with MLLMs. To resolve the need for large-scale region-text training data, we carefully design a pseudo-labeling pipeline for visually-enriched and spatially-localized captions. Our pre-trained encoder is essentially a drop-in replacement of CLIP, with competitive image-text performance, and extra capability demonstrated in region-text tasks and VQA tasks with MLLMs.
Acknowledgment
@@
Appendix A Reproducibility Statement
We made our best efforts to exhaustively state the implementation details. Training hyper-parameters and model architectures are discussed in Section 3.23.3, and 5.1, with a summary in Appendix B and Table A. For evaluation, as mentioned in Section 5.1, we strictly follow the official setup with the codebase released by the original authors if applicable, with details provided in Section B.2. For our datasets, we provide data processing details in Section 4 and example codes in Appendix C. We are working hard on releasing the annotations with internal approvals.
Appendix B Experiment Details
We provide the omitted experiment details for pre-training and the downstream evaluation tasks.
B.1 Pre-training Hyper-parameters
For pre-training both the in-house CLIP baseline and CLOC, we mainly follow the hyper-parameters in (Radford et al., 2021) to train on our in-house datasets. The training images are identical for CLIP and CLOC, while CLOC is trained on the extra region-text annotations of the same images via the proposed VESL pipeline (details in Section C). Table A summarizes the general training hyper-parameters used for all experiments and the setup for components specific to CLOC.
@@ B.1 Pre-training Hyper-parameters
B.2 Evaluation Tasks
We provide more details about the tasks constructed for evaluating the encoders in Section 5.
@@ B.2 Evaluation Tasks
Our CLOC training augments a new capability for CLIP to generate region-level embeddings. This enables us to perform zero-shot region-text tasks, in analogy to the image-text zero-shot tasks like ImageNet classification and COCO text-image retrieval that CLIP has been evaluated on.
@@
For region retrieval, similarly, the CLOC model encodes both the image regions and the region captions from the public region-text GRIT dataset that the regions are annotated by the Kosmos-2 pipeline (Peng et al., 2023). We randomly sampled a $2$K image validation set for fast evaluation. We have verified it is statistically stable compared to the whole set that contains about $20$M in total. Unlike image-text retrieval the image captions are likely unique, the objects in regions of many images might be duplicated. Therefore, we opt to report recall@10 rather than recall@1 for GRIT region retrieval in Table 2.
@@
To demonstrate our CLOC can benefit MLLM end tasks as a better image backbone, we consider two sets of MLLM experiments.
First, we experiment with the Ferret MLLM that is capable of taking spatial referring inputs for grounding and referring VQA tasks [^2] We use the official Ferret codebase: https://github.com/apple/ml-ferret.. Ferret can consume a point, a bounding box, or a free-form referring. It designs a quite complicated visual sampler module that involves point sampling and $k$NN grouping. We suggest the readers refer to Figure 3 and Section 3.2 in (You et al., 2023) for more details. Here we consider two variants of use cases of CLOC compatible with Ferret: (1) we only take the ViT encoder in CLOC to replace the CLIP ViT and still use the original Ferret visual sampler or (2) we further replace the visual sampler with our simple Prompter (essentially a lightweight transformer encoder with box positional encodings) in Section 3.3 as illustrated in Figure 1(3b). More specifically, we simply convert all types of spatial referring as boxes. As evidenced by Table 3 and Table 4, our Prompter can indeed be a much simpler alternative and may perform even better as it is more consistent with CLOC pre-training.
Second, we evaluate on general VQA tasks that do not consider extra spatial referring inputs. The pre-trained ViT of CLOC is a drop-in-replacement of CLIP ViT in two sets of experiments of LLaVA-1.5 (Liu et al., 2023) and LLaVA-NeXT (Liu et al., 2024). The main difference includes different supervised fine-tuning (SFT) sets. Also, LLaVA-NeXT uses the AnyRes technique that decomposes an image into several subimages that are encoded independently with the ViT and concatenated together as the input for the decoder. LLaVA-1.5 by default freezes the ViT while LLaVA-NeXT fine-tunes all parameters during SFT. Since the official LLaVA-NeXT is trained on some proprietary datasets that are not reproducible, we use the Open-LLaVA-NeXT repository [^3] https://github.com/xiaoachen98/Open-LLaVA-NeXT. Our experiments in Table 5 demonstrate CLOC not only slightly improves such general VQA besides Ferret tasks but also generalizes well for both LLaVA-1.5 and LLaVA-NeXT settings.
Appendix C VESL Data Engine
We provide more information about our pseudo-labeling data pipeline proposed in Section 4.
C.1 Implementation Details
@@
For the re-captioning, the goal is to replace AltText with long, diverse, and detailed captions that can be used to generate more visual concepts as the region candidate phrases for the OV detector. Technically, any strong image captioner can be an option. In our paper, we adopt the VeCap pipeline (Lai et al., 2024) and leverage their images with enriched captions.
To extract region phrase candidates from the long captions, we adopt name entity recognition (NER) to extract leaf entities from the captions, inspired by (Zhang et al., 2022). The code listing below shows the Python example implementation, where stop-words and common generic words are filtered, following (Minderer et al., 2024).
Generating bounding boxes and assigning region captions can be done by querying an OV objection detector. We adopted the OWLv2 detector (Minderer et al., 2024) with their pre-trained L/14 checkpoint [^4] OWLv2 CLIP L/14 ST+FT in: https://github.com/google-research/scenic/tree/main/scenic/projects/owl_vit to annotate inputs with $448\times 448$ image resolutions.
@@
Listing 1: Python example codes for Step 2 of VESL in Section 4 for extracting text candidate queries from a caption.
C.2 More Visualizations
As mentioned in the remarks of Section 4, we found the AltText sourced from the original web-crawled images might not have enough details describing the subimage content, thus limiting the diversity and quality of the text candidate queries for the OV detector to detect more meaningful objects. In Figure A we show some cherry-picked examples (since the web-crawled images are quite noisy) just to demonstrate the reasons why high-quality captions can help our region-text annotation pipeline. In (Minderer et al., 2024), the queries are generated by the $n$-grams of the AltText, while ours are by NER as described in Section C.1 on top of the visually-enriched re-captions. Note that, in both methods we use the same pre-trained OV detector but with different approaches to generate the queries.
As shown in Figure A, for easier images like the first row, both methods are doing reasonably well to detect “message card”. However, when the scene becomes complicated (e.g., the second row), our methods can detect more objects since more visual concepts can be extracted from our rich caption as queries for the detector. Similarly, it can be seen that our method captures more items that the AltText missed, e.g., “banana”, “eggs”, “butter”, etc in the third row; “drawstring” in the fourth row; “apples” and “vases” in the last row. Also, it is more likely to extract a more detailed description of the region rather than a class name, such as “green-roofed cottage nestles” in Figure 3 and “decorative metal tree sculptures” in the image in the last row of  Figure A. We believe such high-quality region labels essentially contribute to better supervision for CLOC pre-training.
Figure A: Examples comparing our VESL and the labeling approach in (Minderer et al., 2024) that directly uses the $n$-grams of the crawled AltText. For VESL, each image is annotated with the visual-enriched caption to replace the AltText, which is used to generate region text candidates that capture the image content better.
Appendix D More Discussions
@@
One limitation for CLOC is the labeling efforts in preparing the training data. As we discussed in Section 1, there are no public large-scale region-text datasets since it is expensive to infer such labels up to the scales we consider here. Unlike previous work (Zhong et al., 2022) that cropping boxes from images for annotating, our VESL inference in image-level thus the cost does not scale with the number of detected regions. With that being said, such inference still requires hundreds of GPUs running in parallel for days to scale up to billions of images. We are working on releasing the annotations to accelerate future research for the community.
For CLOC, we focus on the training objective and framework formulation, while making minimal efforts on hyper-parameter tuning, architecture search, dataset cleaning, and etc., thus better performance could be achieved. Besides, although we have included extensive standard evaluation tasks, the fine-grained region knowledge could also be useful on more other under-explored tasks.
@@
We suggest promising future directions. In Section 3.2, our Prompter formulation can take flexible prompts to guide the embeddings for specific tasks. In this work, we consider a prompt as a single bounding box or a text caption, but it has the potential to expand to various types such as points, a mask, users’ free-form referring, or multiple prompts in multiple types together. We think a more versatile Prompter with co-designs for different objectives can have a big potential. Similarly, our VESL labeling pipeline limits to detection box format. Annotators supported for more formats may further boost it. We believe our approach is promising, as more attention has been drawn recently for better re-captions (Li et al., 2024; Fan et al., 2024) that VESL relies on. In addition, CLOC model provides a new capability to extract region features without further training, and thus can be used as a foundation model for exploring new VL applications.
@@
We comment on the computation cost of our framework. Our large models (ViT L/14) were trained on $1024$ v5p TPUs for about $6$ days. To optimize Equation 2, CLOC needs extra computation. The main overheads come from the contrastive matrix but not the lightweight Prompter. Fortunately, we found it feasible since (1) only a few boxes in each image need to be sampled per update; (2) the loss computation becomes a smaller proportion when the ViT scales up. Overall, we found the computation acceptable compared to CLIP. More memory-efficient optimization like SigLIP (Zhai et al., 2023) can be implemented with JAX shard_map [^5] https://jax.readthedocs.io/en/latest/jep/14273-shard-map.html ops.
Discussions on design rationals.
@@
(1) Why not use a local-enhanced encoder? We would like to note that many encoders with great localization like DINOv2 (Oquab et al., 2023), OWLv2 (Minderer et al., 2024), CLIP-Self (Wu et al., 2023), etc. are developed specifically for dense vision tasks that cannot perform image zero-shot tasks like CLIP and CLOC. We would like to emphasize that our goal is to build a drop-in-replacement of CLIP encoder with better localization, without sacrificing CLIP’s original capabilities such as image zero-shot tasks and its important backbone position for MLLMs. Furthermore, perhaps well-known within the MLLMs community, these encoders have been shown in recent reports that they are not comparable enough to compete with CLIP as the vision backbone for MLLM tasks (Tong et al., 2024) due to CLIP’s superiority in vision-language alignment. We thus believe enhancing CLIP itself is more demanding as this paper focuses on.
@@
Although it sounds like a plausible approach, we would like to point out that contrastive pre-training and object detection are fundamentally quite different in their technical rationales. CLIP pre-training is often on large batches of low-resolution and noisy images, while object detection is trained on small batches of high-resolution images. CLIP is by default trained from scratch and object detection is typically initialized from pre-trained encoders and focuses on the detection head. Furthermore, detection requires heavy computation on box proposals to detect all boxes appearing in an image, while our region-text contrastive design allows us to flexibly sample fewer regions per image as motivated in Equation 3. Overall, their data pipeline and distributed training setup are not on the same scale thus such joint training may not be very reasonable.
With that being said, some previous works do have attempts that are the exceptions but only for some but not all of the mentioned aspects, and mainly for the purpose of detection. For instance, DetCLIP-v2 (Yao et al., 2023b) adds image-text contrastive loss into detection loss to improve open-vocabulary capability for detection. OWLv2 pre-trains the detector with rather small resolutions but still with a batch size of a maximum $256$ since each image will need to predict up to $100$ boxes during training. Both DetCLIP-v2 and OWLv2 fine-tune from a pre-trained encoder.
On the contrary, we study pre-training the encoder from scratch, which may be complementary to the previous efforts. CLOC maximizes the similarity in co-design with CLIP, thus making it much easier to develop within the same codebase.
(3) Do we really need to train CLOC from scratch? What if we fine-tune from CLIP? As CLIP pre-training is expensive, one may wonder if it is necessary to train from scratch on the proposed region-text datasets, or if we can initialize from a standard CLIP trained on image-text pairs only and fine-tunes with CLOC for a shorter stage. Our early investigation, even with extensive hyper-parameter tuning, suggests it is likely to be suboptimal compared to training from scratch directly. For instance, we initialize from the CLIP model 2 in Table 2 and fine-tunes it for another extra $100$K steps with the CLOC training loss Equation 4. The model reaches $64.1\%/19.1\%$ mAcc on COCO/LVIS region recognition, which is much worse than $70.1\%/27.2\%$ of the trained-from-scratch model 8, even with more overall training steps.